# Chaotic discretization theorems for forced linear and nonlinear coupled oscillators

### v1

@@ 5.2 Bifurcation analysis
Given Proposition 3, we perform the bifurcation diagram of (2) by taking (51). In particular, we construct two separate lists $y_{even}$, $y_{odd}$ for the even and odd iterations respectively; then, we alternatively plot the values of $y_{even}$ and $y_{odd}$ in order to obtain the whole orbit. From a computationally point of view, the previous method is neither more nor less expensive than dealing with the four-dimensional system; however, we obtain slightly more clear results.
@@
Remark 3.
Discontinuities exhibited in Figure 9 are not totally unexpected. Since the map (2) presents discontinuities of the first kind at $y_{n}=2k\pi$, $k\in\mathbb{Z}$, we are supposed to see analogous jumps in the bifurcation diagram for $y_{n}$.
In Figure 10 - 12 they are shown magnifications of the parts of interest in Figure 9.
Figure 10: Magnification of Figure 9 for the upper logistic-like subdiagram.
Figure 11: Magnification of Figure 9 for the lower logistic-like subdiagram.
Figure 12: Magnification of Figure 9 for the chaotic regime on the right.
The system exhibits the typical period doubling cascade, encountered in the well known logistic map May as well as in several discrete maps; however, it arises in a peculiar way.
@@
    The system starts with periodic orbits of period $3$, 66, 1212 up to $\gamma\sim 0.1$.
@@
Now, we perform a bifurcation analysis of the system (2) in order to confirm the period doubling bifurcations shown in Figure 9. We use the software MatContM, a MATLAB tool devoted to bifurcation analysis of iterated smooth maps through continuation methods; see Dhooge for a review of the native software MatCont (developed for ODE systems) and Neirynck for an application of MatContM to a nonlinear map of economic interest.
As before, we implement the one-dimensional map (51); in particular, we consider the even orbit by setting $\tilde{\alpha}=-\alpha=-1$, $\beta=0.1$ and take $\gamma$ as bifurcation parameter. In Figure 13 it is shown the continuation diagram for fixed points and period doubling bifurcations. We find period doubling bifurcations at $\gamma=0.10516472$ and at $\gamma=2.5912211$, according to the bifurcation diagram in Figure 9. It is also found a generalized period doubling bifurcation Kuznetsov at $\gamma=0.52673498$.
@@
We conclude this section by presenting in Figure 14 Lyapunov exponents for the system (2) for $\alpha=-\tilde{\alpha}=-1$, $\beta=0.1$, $\gamma\in[0,5]$; we implement the even orbit for (51) with initial condition $y_{0}=0.1$. Negative Lyapunov exponents are found for values of $\gamma$, according to the stability islands shown in Figure 9 - 12.
Figure 14: Lyapunov exponents of the map (2) for the variable $y_{n}$ with respect to $\gamma\in[0,5]$ and for $\tilde{\alpha}=-\alpha=1$, $\beta=0.1$, $y_{0}=0.1$.
@@ 6 Conclusions
Chaotic behavior of the modular definition of the map has been numerically confirmed by identifying a strange attractor for general choices on parameters and initial conditions, showing the sensitive dependence of the system on initial data and the presence of positive Lyapunov exponents. Furthermore, the bifurcation diagram of the map exhibits a fractal structure analogous to the logistic map, with a singular rising of the period doubling cascade.
@@
The rising of a strange attractor for the system leads to some questions. Firstly, we might numerically compute the Hausdorff dimension Hausdorff of the attractor, in order to have a confirmation of its fractal dimension. Secondly, it is not immediate to provide an explanation for certain dynamical properties, such as the rising of the attractor only on certain projection planes or the peculiar regular orbits rising in the symmetric case. Finally, a deeper bifurcation analysis of the system may be performed, in order to find further bifurcation points and provide a complete scenario for stability and transition to chaos for the system.
@@ Acknowledgments
The first author is grateful to prof. Davide Liessi (University of Udine) for his insightful lectures on the MATLAB tool MatCont. The authors also thank prof. Michele Miranda (University of Ferrara) for helpful discussions regarding the initial flawed proof of Theorem 3.

### v2

@@ 5.2 Bifurcation diagrams and Lyapunov exponents
Given Proposition 4, we begin performing the bifurcation diagram of (56). We implement the classical algorithm Strogatz [2015] for the construction of the bifurcation diagram. The basic idea of the algorithm is to compute the trajectory of the system for every value of the bifurcation parameter up to a certain iteration step ($n_{it}$), discarding the first iterations ($n_{tran}$) so that the system has actually reached its attractor and any transitory effects are not shown. In particular, we construct two separate lists $y_{even}$, $y_{odd}$ for the even and odd iterations respectively; then, we alternatively plot the values of $y_{even}$ and $y_{odd}$ in order to obtain the whole orbit.
@@
Figure 10: Magnification of Figure 9 related to the blue rectangle.
Figure 11: Magnification of Figure 9 related to the green rectangle.
Figure 12: Magnification of Figure 9 related to the red rectangle.
The system exhibits the typical period doubling cascade, encountered in the well known logistic map May [1976] as well as in several discrete maps; however, it arises in a peculiar way.
@@
    The system starts with periodic orbits of period $3$, 66, $12$ up to $\gamma\sim 0.1$.
@@
Discontinuities exhibited in Figure 9 are not totally unexpected. Since the map (56), such as (7), presents discontinuities of the first kind at $y_{n}=2k\pi$, $k\in\mathbb{Z}$, we are supposed to see analogous jumps in the bifurcation diagram for $y_{n}$.
In discontinuous systems, multistability is a frequently observed phenomenon Pisarchik and Feudel [2014], defined as the coexistence of different attractor and completely different behavior for a given choice on the parameters, depending solely on the initial conditions. To investigate the existence of multistable states, in Figure 13 we present basins of attraction of (7) for values of the parameters corresponding to the second discontinuity shown in Figure 9. The simulations were performed using the Julia scripts provided in Wagemakers [2025], designed for generating basins of attraction for iterated maps and systems of ordinary differential equations. Figure 13 shown the projection of the basins of attraction onto the $(x,y)$ plane for $\alpha=-\tilde{\alpha}=-1$, $\beta=0.1$, $\gamma=3.3788$, with fixed $z_{0}=\omega_{0}=0$ and $x_{0},y_{0}\in[0,2\pi)$. The plot clearly shows the coexistence of at least three distinct attractors, identified by different colors corresponding to the asymptotic state reached by the initial conditions, confirming the presence of multistability in the system (7); furthermore, the basins exhibit a clear fractal structure.
Figure 13: Basins of attraction of (7), projection onto the $(x,y)$ plane. Parameters: $\alpha=-\tilde{\alpha}=-1$, $\beta=0.1$, $\gamma=3.3788$. Fixed initial conditions: $z_{0}=\omega_{0}=0$. White, red and black regions correspond to distinct asymptotic states reached for a given set of initial condition $(x_{0},y_{0})$.
Now, we perform a bifurcation analysis of the system (56) in order to confirm the period doubling bifurcations shown in Figure 9. We use the software MatContM, a MATLAB tool devoted to bifurcation analysis of iterated smooth maps through continuation methods; see Dhooge et al. [2008] for a review of the native software MatCont (developed for ODE systems) and Neirynck et al. [2016] for an application of MatContM to a nonlinear map of economic interest.
As before, we implement the even orbit of (56), set $\tilde{\alpha}=-\alpha=-1$, $\beta=0.1$ and take $\gamma$ as bifurcation parameter. In Figure 14 it is shown the continuation diagram for fixed points and period doubling bifurcations. We find period doubling bifurcations at $\gamma=0.10516472$ and at $\gamma=2.5912211$, according to the bifurcation diagram in Figure 9. It is also found a generalized period doubling bifurcation Kuznetsov [2023] at $\gamma=0.52673498$.
@@
In Figure 15 we present Lyapunov exponents for the system (56) for $\alpha=-\tilde{\alpha}=-1$, $\beta=0.1$, $\gamma\in[0,5]$; we implement the even orbit for (56) with initial condition $y_{0}=0.1$. Negative Lyapunov exponents are found for values of $\gamma$, according to the stability islands shown in Figure 9 - 12.
Figure 15: Lyapunov exponents of the map (56) for $y_{0}=0.1$ with respect to $\gamma\in[0,5]$. Parameters: $\tilde{\alpha}=-\alpha=1$, $\beta=0.1$, $y_{0}=0.1$.
Numerical results presented in Figure 9 - 12, 14 - 15 refer to the 1D map (56), whose dynamics is equivalent to the four-dimensional (7) when $\alpha=-\tilde{\alpha}$. Now, we present 2D bifurcation diagrams for the full system (7) and spectra of Lyapunov exponents, that were generated using a dedicated Python implementation. For a comprehensive overview of a recent high-performance numerical tool for bifurcation analysis, we refer to Rybin et al. [2024].
We perform bifurcation analysis for three sets of parameters. For all the following simulations we consider $n_{it}=5.000$ iterations by discarding the first $n_{tran}=500$ ones and discretizing the parameters range into $1.000$ values. The color bar indicates the maximum Lyapunov exponents of (7) for a given pair of parameters.
In Figure 16 it is shown the 2D bifurcation diagram $(\beta,\gamma)$ for $\alpha=-\tilde{\alpha}=-1$, $\beta\in[0,5]$, $\gamma\in[0,5]$. In Figure 17 the Lyapunov spectrum for same values of $\alpha$, $\tilde{\alpha}$ and $\beta=0.1$ is presented, by varying $\gamma$ in the interval $[0,5]$.
Figure 16: 2D bifurcation diagram $(\beta,\gamma)$ for the map (7). Parameters: $\alpha=-\tilde{\alpha}=-1$.
Figure 17: Lyapunov exponents spectrum for the map (7). Parameters: $\alpha=-\tilde{\alpha}=-1$, $\beta=0.1$.
In Figure 18 it is shown the 2D bifurcation diagram $(\alpha,\gamma)$ for $\tilde{\alpha}=2$, $\beta=0.8$, $\alpha\in[-5,0]$, $\gamma\in[0,10]$. In Figure 19 the Lyapunov spectrum for same values of $\tilde{\alpha}$, $\beta$ and $\alpha=-2.5$ is presented, by varying $\gamma$ in the interval $[0,10]$. These results are comprehensive of the parameters choice for the confined strange attractor shown in Figure 3 - 4.
Figure 18: 2D bifurcation diagram $(\alpha,\gamma)$ for the map (7). Parameters: $\tilde{\alpha}=2$, $\beta=0.8$.
Figure 19: Lyapunov exponents spectrum for the map (7). Parameters: $\tilde{\alpha}=2$, $\beta=0.8$, $\alpha=-2.5$.
In Figure 20 it is shown the 2D bifurcation diagram $(\alpha,\beta)$ for $\tilde{\alpha}=2$, $\gamma=4$, $\alpha\in[-1,1]$, $\beta\in[-1,1]$. In Figure 21 the Lyapunov spectrum for same values of $\tilde{\alpha}$, $\gamma$ and $\alpha=0$ is presented, by varying $\beta$ in the interval [1,1][-1,1].
Figure 20: 2D bifurcation diagram $(\alpha,\beta)$ for the map (7). Parameters: $\tilde{\alpha}=2$, $\gamma=4$.
Figure 21: Lyapunov exponents spectrumfor the map (7). Parameters: $\tilde{\alpha}=2$, $\gamma=4$, $\alpha=0$.
A few remarks are in order. Firstly, notice that the Jacobian of (7) exhibits a symmetry on its eigenvalues such that Lyapunov exponents are equal in pairs, as evident from the spectra shown in Figure 17, 19, 21. Furthermore, for the parameters set selected in Figure 19, we have a positive sum of the Lyapunov exponents, that is incompatible with a global strange attractor; however, a confined strange attractor can still emerge in specific projections, such as the one that is observed in Figure 3 - 4. Finally, for all the previous simulations wide sets of parameters exhibit two positive Lyapunov exponents, leading to hyperchaos Kapitaniak et al. [2000], Stankevich et al. [2020] for the system (7).
@@ 6 Conclusions
Chaotic behavior of the modular definition of the map has been numerically confirmed by identifying a confined strange attractor for general choices on parameters and initial conditions. The one-dimensional bifurcation diagram of the map exhibits a fractal structure analogous to the logistic map, with a singular rising of the period doubling cascade. Sensitive dependence of the system on initial data and rising of positive Lyapunov exponents are confirmed by two-dimensional bifurcation diagrams and the Lyapunov exponents spectra. Furthermore, the basins of attraction of the system confirm the presence of multistable states.
@@
The rising of a strange attractor for the system leads to some questions. Firstly, we might numerically compute the Hausdorff dimension Hausdorff [1919] of the attractor, in order to have a confirmation of its fractal dimension. Secondly, it is not immediate to provide an explanation for certain dynamical properties, such as the rising of the attractor only on certain projection planes or the peculiar regular orbits rising in the symmetric case. Finally, a deeper bifurcation analysis of the system may be performed, in order to find further bifurcation points and provide a complete scenario for stability and transition to chaos for the system.
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author used Google Gemini (Large Language Model) in order to write the Python code for reproducing the 2D bifurcation diagrams and the Lyapunov exponents spectra (Figure 16 - 21). After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Acknowledgments
The first author is grateful to prof. Davide Liessi (University of Udine) for his insightful lectures on the MATLAB tool MatCont. The authors also thank prof. Michele Miranda (University of Ferrara) for helpful discussions regarding the initial flawed proof of Theorem 3. The authors are grateful to the anonymous reviewers for their helpful comments and suggestions.